# Characterizing Open-Ended Evolution Through Undecidability Mechanisms in Random Boolean Networks

### v1

@@ Discussion
Updating York’s OEE classification, a group in Tokyo refined such a taxonomy (packard2019overview). Our metric most directly instantiates Tokyo Type 1 OEE (ongoing generation of interesting new kinds of entities/interactions) when the phenotype map treats attractors or coarse-grained behaviors as entities; it can also serve as an indicator for the evolution of evolvability (Tokyo Type 2 OEE) when the rate of novelty production increases across windows or under meta-evolution of the genotype-phenotype map. Where a system undergoes coarse-graining shifts that produce qualitatively new levels of organization (major transitions; Tokyo Type 3 OEE), the resulting explosion of new attractor classes will register as marked increases in $\Omega$. By construction, our metric remains agnostic about meaning; to speak to Tokyo Type 4 OEE (semantic evolution) requires an explicit interpretation function from system states to task- or environment-defined semantics, at which point one can track whether the new phenotypes also carry new semantic roles. In short, $\Omega$ equips us with a portable finite-time detector for the axis of “ongoing generation” that both the York and Tokyo taxonomies place at the center of OEE. Now, let us go back and interpret the results we obtained.
In homogeneous networks, the PBN transformation achieves the largest $\Omega$ at low-medium connectivities, whereas Modal logic dominates at high connectivities (Fig. 1). The PBN outcome is consistent with the mechanism: piecewise deterministic motion inside a context, punctuated by rare switches, tends to string together long residence times in distinct attractor families, inflating the residence-time-weighted cycle contribution in $\Omega$. Biologically, this mirrors bet-hedging via phenotype switching under resource-scarce conditions: exactly where sparse coupling (low $K$) limits within-context exploration, so context changes do the heavy lifting (acar2008stochastic).
By contrast, when coupling is dense (large $K$), Modal gating becomes advantageous. “Necessary” constraints act as context-dependent canalizing conditions that suppress runaway sensitivity, while “possible” gates open alternative routes when local context permits; together they tame chaotic overshoot and create metastable corridors that repeatedly access distinct limit cycles. As a practical rule: in synchronous, homogeneous systems with limited connectivity, stochastic program switching (PBN) yields the highest open-endedness; as connectivity grows, context-dependent gating (Modal) overtakes by structuring exploration without freezing it.
Moving to heterogeneous networks, the picture changes qualitatively (Fig. 2). The PBN curve collapses toward zero over the entire $K$ range, while Modal again excels at higher $K$, and Paraconsistent and Quantum-inspired mechanisms show localized gains at low-intermediate $K$. This collapse is not due to a change in the number of point attractors, those are determined by the rule tables and are independent of the updating scheme (Gershenson2004b), but rather to how our stochastic asynchronous updates interact with structural heterogeneity. In our simulations we use CUBEWALKERS’ asynchronous_set update scheme, which randomly selects a set of nodes to be updated at each time step, so that the dynamics along a single trajectory are effectively noisy and long synchronous cycles are rarely realized. Combined with an Exponential in-degree distribution, this drives most initial conditions quickly into small recurrent classes (often fixed points or very short cycles) that have large basins of attraction.
Within each PBN context, trajectories therefore spend almost all their time in short attractors with cycle lengths $k_{j}$ of order one; context switches mostly move the system between the basins of these same few attractors, rather than uncovering new long cycles. As a result, the summed residence-cycle contribution $\sum_{j}d_{j}k_{j}$ grows at most linearly with $T$, so $\Omega\rightarrow 0$ despite ongoing context changes. Modal logic is the exception because its necessity/possibility tests provide conditional activation that intermittently re-opens pathways even under asynchronous settling; hence its large spikes at high $K$. Paraconsistent logic yields a broad shoulder around $K\approx 2.3$: introducing controlled local contradictions near the order-critical boundary sustains exploratory dynamics without exploding to noise, in line with the known criticality-extending effects of heterogeneity (sanchez2023heterogeneity; lopez2023temporal). Quantum-inspired rules perform best at very low $K$: superposition tokens plus a few long-range entanglements furnish correlated branching when wiring is sparse, boosting early-time exploration before asynchronous contraction dominates. The strategy that follows is therefore different from the homogeneous case: under heterogeneity/asynchrony, prefer Modal gating at high connectivity and use Paraconsistent or Quantum-inspired rules to lift $\Omega$ in sparse networks.
Some patterns are trivial yet worth stating explicitly. In both regimes, the Classical curve sinks over the entire $K$ range because a deterministic network visiting a single attractor contributes $\Omega\rightarrow 0$ for large enough values of $T$ (see Supplementary Material); $\Omega$ rewards persistent traversal of multiple cyclic phenotypes, not mere residence in one. The consistently poor performance of ARM is also instructive: per-read truth-table mutation anneals away stable local structure, pushing dynamics toward effectively memoryless transitions; this drives down the product of cycle length and dwell time and is qualitatively consistent with the classic annealed-approximation intuition that high flip rates erase organized attractor geometry (derrida1986random). In short, ARM provides plasticity but not the kind of structured, state-dependent plasticity that $\Omega$ rewards.
@@ Conclusion
Our results allow us to propose the following actionable design principles for maximizing open-endedness:
@@
    For heterogeneous/asynchronous systems: Modal at high $K$ and Paraconsistent or Quantum-inspired logic at low-intermediate $K$ maximize open-endedness.
@@
We note two cross-checks that bolster these interpretations. First, the high-$\Omega$ zones for Paraconsistent and Modal variants align with regions where heterogeneity is known to extend or reshape the critical regime in Boolean networks (lopez2023temporal). Second, the PBN advantage under homogeneity but not heterogeneity is consistent with the literature on update schemes: when all nodes are updated synchronously, the dynamics follow a fixed deterministic map that can sustain rich cyclic structure, whereas stochastic asynchronous updates tend to break long cycles and contract dynamics into small recurrent classes (Gershenson2004b). In our heterogeneous simulations we combine an Exponential in-degree distribution with CUBEWALKERS’ non-deterministic asynchronous_set scheme, so it is this added randomness in the update order, rather than asynchronicity per se, that erodes the attractor geometry that $\Omega$ rewards. Deterministic asynchronous schemes with fixed node orderings can, by contrast, preserve or even lengthen cycles, so our claim is restricted to the non-deterministic case we study here.
@@
In the aggregate, Figures 1-2 suggest that state-dependence is necessary but not sufficient; how it is introduced—via context switching, conditional gating, controlled contradictions, or correlated branching—interacts strongly with architectural conditions (connectivity and update heterogeneity) to determine whether open-ended dynamics flourish or fade. This points towards a design principle for engineering OEE: choose the mechanism whose mode of uncertainty (what is randomized, when, and where) best complements the system’s wiring and timing. In this sense, our results are analogous to the design of the right amount of antifragility in Boolean networks—pineda2019novel showed that the benefit a system derives from perturbations depends on how the level and pattern of external noise are matched to the underlying dynamics, and identified regimes where perturbations improve a satisfaction functional rather than merely degrading performance. Here we obtain a complementary view: instead of tuning environmental noise for a fixed network, we tune the internal realization of uncertainty for a given architecture, and assess its effect on the long-term production of novel attractors via $\Omega$. Still, our approach has focused specifically on discrete state spaces, leaving aside systems with continuous (or hybrid) state spaces. In the following, we sketch a principled extension of our metric that preserves its core properties while making it applicable beyond purely discrete models.
Let $(X,d)$ be a compact metric space and $\Phi_{t}$ a measurable flow (or map) on $X$. Choose (i) a finite measurable partition $\Pi=\{C_{1},\dots,C_{Q}\}$ of $X$ that represents the phenotypic coarse-graining of interest and (ii) a tolerance $\varepsilon>0$ that encodes measurement noise or resolution. Given a trajectory $x(t)$, form the symbolic stream $s(t)\in\{1,\dots,Q\}$ by setting $s(t)=q$ whenever $x(t)\in C_{q}$. We say that the dynamics enters an $(\varepsilon,\Pi)$-attractor episode $A_{j}^{\varepsilon,\Pi}$ at time $t_{j}^{\mathrm{in}}$ if thereafter the orbit remains in a trapping region $\mathcal{T}\subseteq X$ and returns within distance $\varepsilon$ of some previously visited state in $\mathcal{T}$ before time $t_{j}^{\mathrm{out}}$; the residence time is $d_{j}^{\varepsilon,\Pi}=t_{j}^{\mathrm{out}}-t_{j}^{\mathrm{in}}$. For that episode define the symbolic cycle-length
defining open-endedness for continuous state spaces as
This extension preserves the key behavior of our discrete metric: (i) if the system converges to a single fixed point or limit cycle contained within one recurrent class of $\Pi$, then $\Omega_{\varepsilon,\Pi}=0$ because there is only one episode and $d_{1}^{\varepsilon,\Pi}=O(T)$ while $K_{1}^{\varepsilon,\Pi}$ is $O(1)$; (ii) if the number of distinct $(\varepsilon,\Pi)$-attractor episodes diverges, then $\Omega_{\varepsilon,\Pi}$ diverges as well. Moreover, by the same Cauchy-Schwarz argument used in the discrete case we obtain a resolution-dependent upper bound
Second, for long trajectories, exact recurrence detection is costly. In continuous state spaces many trajectories will cross separatrices, escape to infinity, or otherwise hit “points of no return”, so exact state revisits are not expected globally. In practice, we would first identify coarse-grained recurrent classes (quasi-attractors) by looking for approximate returns to the same cell of a finite-resolution partition. Our GPU streaming/extractor could be extended to use fast $\varepsilon$-recurrence queries (e.g., locality-sensitive hashing of Poincaré samples or recurrence-plot-based triggers), guaranteeing $O(1)$ amortized checks per step. Trajectories that never return to a cell at the chosen resolution simply do not generate a recurrent class there, and thus contribute $\Omega_{\varepsilon,\Pi}=0$ at that scale. Once such coarse-grained recurrent classes have been detected, their residence times $d_{j}^{\varepsilon,\Pi}$ become random variables. We then estimate $\Omega_{\varepsilon,\Pi}$ by ergodic averages over long runs, relating it to return-time statistics within each coarse-grained recurrent class (i.e., quasi-attractor), rather than assuming global recurrence of the full system.
@@
Since we just came up with a continuous generalization of our metric, we could apply it to continuous GRN models (such as Hill-function ODEs (ehsan2018method), piecewise-linear Glass systems (edwards2007periodicity), random fuzzy networks (Zapata2020), etc.) using phenotype-aware partitions and compare with the discrete-RBN results to test which undecidability-adjacent mechanisms (modal gating, controlled contradictions, context switching) remain most effective in the continuous limit. Moreover, we could relate $\Omega_{\varepsilon,\Pi}$ to invariants such as topological entropy (adler1965topological) or return-time spectra in chaotic attractors (boev2014poincare) and determine the conditions under which $\lim_{\varepsilon\to 0}\lim_{\Pi\to\text{refined}}\Omega_{\varepsilon,\Pi}$ exists and is equal to the discrete-state $\Omega$ for Boolean embeddings. This opens a path to compare, on equal footing, discrete and continuous regulatory schemes across modeling paradigms, and to design interventions that promote sustained novelty under real biophysical constraints.
Acknowledgements
The authors will acknowledge the anonymous reviewers who will enrich our contribution with their valuable comments.
@@ Acknowledgements
The code used to perform the simulations in this study can be found at https://github.com/amahury/OEE-metric (accessed on 16 December 2025).
@@
C.G. acknowledges support from the SSIE School.

### v2

@@ Discussion
Updating York’s OEE classification, a group in Tokyo refined such a taxonomy (Packard et al., 2019a). Our metric most directly instantiates Tokyo Type 1 OEE (ongoing generation of interesting new kinds of entities/interactions) when the phenotype map treats attractors or coarse-grained behaviors as entities; it can also serve as an indicator for the evolution of evolvability (Tokyo Type 2 OEE) when the rate of novelty production increases across windows or under meta-evolution of the genotype-phenotype map. Where a system undergoes coarse-graining shifts that produce qualitatively new levels of organization (major transitions; Tokyo Type 3 OEE), the resulting explosion of new attractor classes will register as marked increases in $\Omega$. By construction, our metric remains agnostic about meaning; to speak of Tokyo Type 4 OEE (semantic evolution) requires an explicit interpretation function from system states to task- or environment-defined semantics, at which point one can track whether the new phenotypes also carry new semantic roles. In short, $\Omega$ equips us with a portable finite-time detector for the axis of “ongoing generation” that both the York and Tokyo taxonomies place at the center of OEE.
To generalize $\Omega$ to continuous (or hybrid) state spaces, let $(X,d)$ be a compact metric space and $\Phi_{t}$ a measurable flow (or map) on $X$. Choose (i) a finite measurable partition $\Pi=\{C_{1},\dots,C_{Q}\}$ of $X$ that represents the phenotypic coarse-graining of interest and (ii) a tolerance $\varepsilon>0$ that encodes measurement noise or resolution. Given a trajectory $x(t)$, form the symbolic stream $s(t)\in\{1,\dots,Q\}$ by setting $s(t)=q$ whenever $x(t)\in C_{q}$. We say that the dynamics enters an $(\varepsilon,\Pi)$-attractor episode $A_{j}^{\varepsilon,\Pi}$ at time $t_{j}^{\mathrm{in}}$ if thereafter the orbit remains in a trapping region $\mathcal{T}\subseteq X$ and returns within distance $\varepsilon$ of some previously visited state in $\mathcal{T}$ before time $t_{j}^{\mathrm{out}}$; the residence time is $d_{j}^{\varepsilon,\Pi}=t_{j}^{\mathrm{out}}-t_{j}^{\mathrm{in}}$. For that episode define the symbolic cycle-length
defining open-endedness for continuous state spaces as
This extension preserves the key behavior of our discrete metric: (a) if the system converges to a single fixed point or limit cycle contained within one recurrent class of $\Pi$, then $\Omega_{\varepsilon,\Pi}=0$ because there is only one episode and $d_{1}^{\varepsilon,\Pi}=O(T)$ while $K_{1}^{\varepsilon,\Pi}$ is $O(1)$; (b) if the number of distinct $(\varepsilon,\Pi)$-attractor episodes diverges, then $\Omega_{\varepsilon,\Pi}$ diverges as well. Moreover, by the same Cauchy-Schwarz argument used in the discrete case we obtain a resolution-dependent upper bound
Second, for long trajectories, exact recurrence detection is costly. In continuous state spaces many trajectories will cross separatrices, escape to infinity, or otherwise hit “points of no return”, so exact state revisits are not expected globally. In practice, we would first identify coarse-grained recurrent classes (quasi-attractors) by looking for approximate returns to the same cell of a finite-resolution partition. Our GPU streaming/extractor could be extended to use fast $\varepsilon$-recurrence queries (e.g., locality-sensitive hashing of Poincaré samples or recurrence-plot-based triggers), guaranteeing $O(1)$ amortized checks per step. Trajectories that never return to a cell at the chosen resolution simply do not generate a recurrent class there, and thus contribute $\Omega_{\varepsilon,\Pi}=0$ at that scale. Once such coarse-grained recurrent classes have been detected, their residence times $d_{j}^{\varepsilon,\Pi}$ become random variables. We then could estimate $\Omega_{\varepsilon,\Pi}$ by ergodic averages over long runs, relating it to return-time statistics within each coarse-grained recurrent class (i.e., quasi-attractor), rather than assuming global recurrence of the full system.
@@ Conclusion
Our experiments use random network ensembles and stylized state-dependent operators as a controlled testbed. We do not claim that specific non-classical logics are literally implemented in gene regulation. Rather, these operators serve as mechanistic motifs for injecting context dependence and structured uncertainty into discrete regulatory dynamics. In empirical applications, $\Omega$ should be computed on curated GRN models with biologically justified update/noise assumptions and interpreted relative to the chosen phenotype coarse-graining and finite simulation horizon. Even so, our results allow us to propose the following actionable design principles for maximizing open-endedness:
@@
    For heterogeneous/asynchronous systems: Modal at high $K$ and Paraconsistent or Quantum-inspired mechanism at low-intermediate $K$ maximize open-endedness.
@@
We note two cross-checks that bolster these interpretations. First, the high-$\Omega$ zones for Paraconsistent and Modal variants align with regions where heterogeneity is known to extend or reshape the critical regime in Boolean networks (López-Díaz et al., 2023). Second, the PBN advantage under homogeneity but not heterogeneity is consistent with the literature on update schemes: when all nodes are updated synchronously, the dynamics follow a fixed deterministic map that can sustain rich cyclic structure, whereas stochastic asynchronous updates tend to break long cycles and contract dynamics into small recurrent classes (Gershenson, 2004b). In our heterogeneous simulations we combine an Exponential in-degree distribution with CUBEWALKERS’ non-deterministic asynchronous_set scheme, so it is this added randomness in the update order, rather than asynchronicity per se, that erodes the attractor geometry that $\Omega$ rewards. Deterministic asynchronous schemes with fixed node orderings can, by contrast, preserve or even lengthen cycles, so our claim is restricted to the non-deterministic case we study here.
@@
In the aggregate, Figures 1-2 suggest that state-dependence is necessary but not sufficient; how it is introduced—via context switching, conditional gating, controlled contradictions, or correlated branching—interacts strongly with architectural conditions (connectivity and update heterogeneity) to determine whether open-ended dynamics flourish or fade. This points towards a design principle for engineering OEE: choose the mechanism whose mode of uncertainty (what is randomized, when, and where) best complements the system’s wiring and timing. In this sense, our results are analogous to the design of the right amount of antifragility in Boolean networks—Pineda et al. (2019) showed that the benefit a system derives from perturbations depends on how the level and pattern of external noise are matched to the underlying dynamics, and identified regimes where perturbations improve a satisfaction functional rather than merely degrading performance. Here we obtain a complementary view: instead of tuning environmental noise for a fixed network, we tune the internal realization of uncertainty for a given architecture, and assess its effect on the long-term production of novel attractors via $\Omega$.
@@
Since we just came up with a continuous generalization of our metric in the Discussion, we could apply it to continuous GRN models (such as Hill-function ODEs (Ehsan Elahi and Hasan, 2018), piecewise-linear Glass systems (Edwards et al., 2007), random fuzzy networks (Zapata et al., 2020), etc.) using phenotype-aware partitions and compare with the discrete-RBN results to test which undecidability-adjacent mechanisms (modal gating, controlled contradictions, context switching) remain most effective in the continuous limit. Moreover, we could relate $\Omega_{\varepsilon,\Pi}$ to invariants such as topological entropy (Adler et al., 1965) or return-time spectra in chaotic attractors (Boev et al., 2014) and determine the conditions under which $\lim_{\varepsilon\to 0}\lim_{\Pi\to\text{refined}}\Omega_{\varepsilon,\Pi}$ exists and is equal to the discrete-state $\Omega$ for Boolean embeddings. This opens a path to compare, on equal footing, discrete and continuous regulatory schemes across modeling paradigms, and to design interventions that promote sustained novelty under real biophysical constraints.
Acknowledgments
The authors acknowledge two anonymous reviewers who have enriched our contribution with their valuable comments and observations.
@@ Acknowledgments
The code used to perform the simulations in this study can be found at https://github.com/amahury/OEE-metric (accessed on 31 March 2026).
Author Contribution
A.J.L.‑D. conceptualized the project, carried out formal analysis, performed simulations, and wrote the original draft. P. J. R. T. provided theory and code related to PBN. G. L. F. developed the extension of the metric to continuous/hybrid state spaces. C.G. acquired funding, provided computational resources, and supervised. All authors validated, reviewed, edited, and wrote the final manuscript.
Conflict of interests
We declare we have no competing interests.
@@
C.G. acknowledges support from the SSIE School, NYS 910247-22.